OpenStudy (anonymous):

Can someone solve this and show work? thanks. Given logb a = x , then bx = a where b > 0 but b #1, and a > 0. 1. 3log2 x = 12 2. log5 125 = x 2. 3 + 4 logx 4 = 5

OpenStudy (anonymous):

\[3\log_{2} X = 12 --- 3+4\log_{x} 4=5 --- \log_{5} 125=x\]

OpenStudy (anonymous):

\[3log_2x = 12 \]Divide both sides by 3:\[log_2x = 4 \implies 2^4 = x = 16\] Now you try the rest.

OpenStudy (anonymous):

And I'll check your answers, or help if you get stuck.

OpenStudy (anonymous):

So the next problem is: \[log_5(125) = x\] Right? How can you re-write that?

OpenStudy (anonymous):

5^x = 125

OpenStudy (anonymous):

That is exactly right. So we need to know what power of 5 is 125. That power will be the x

OpenStudy (anonymous):

So 125 is which power of 5?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Correct. So if: \[5^x = 5^3\] What must x = ?

OpenStudy (anonymous):

3? not sure.

OpenStudy (anonymous):

Yep. 5^x = 5^3 means x = 3